# Bias-Field Digitized Counterdiabatic Quantum Algorithm for Higher-Order Binary Optimization

### v1

@@ IV Conclusion
In this study, we investigate and evaluate the performance of novel variants of the recently proposed BF-DCQO algorithm [48] by solving large-scale HUBO instances on IBM quantum processors. The superiority of our BF-DCQO variant over well-established classical techniques and optimization on D-Wave is verified through computing the approximation ratio and the distance to solution in an iterative process. We obtain better optimal solutions while the circuits after each iteration demand less resources. Furthermore, in order to address not only the current but also the near-term hardware, we examine the expected performance of BF-DCQO on the upcoming IBM Osprey platform with 433 qubits. Simulating 433-qubit HUBO instances without noise yields promising results. The BF-DCQO algorithm does not require any classical optimization subroutines, thus avoiding potential trainability drawbacks, nor does it require extra qubits for mapping the initial HUBO into a QUBO problem. It thus emerges as a suitable algorithm for solving higher-order binary optimization problems on both current and near-term quantum platforms. We believe that with these results it is fair to claim that the commercial quantum advantage era has started.
Acknowledgements.
@@
We thank Anton Simen Albino, Gaurav Dev and Balaganchi A. Bhargava for fruitful discussions. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendix A First-order nested commutator calculation
In order to include the counterdiabatic driving contribution into our protocol, we have to perform the map $H_{\text{ad}}(\lambda)\mapsto H_{\text{ad}}(\lambda)+\dot{\lambda}\alpha_{1}(\lambda)O_{1}(\lambda)$ with $O_{k}=[H_{\text{ad}}(\lambda),O_{k-1}(\lambda)]$ starting from $O_{1}(\lambda)=[H_{\text{ad}}(\lambda),\partial_{\lambda}H_{\text{ad}}(\lambda)]=[H_{i},H_{f}]$ and $\alpha_{1}(\lambda)=-\Gamma_{1}(\lambda)/\Gamma_{2}(\lambda)$ with $\Gamma_{k}(\lambda)=\text{tr}\big{[}O_{k}^{\dagger}(\lambda)O_{k}(\lambda)\big{]}$. We start by computing
$\begin{split}O_{1}(\lambda)&=-2i\sum_{i}h_{i}^{x}h_{i}^{z}\sigma^{y}_{i}-2i\sum_{i<j}J_{ij}(h_{i}^{x}\sigma^{y}_{i}\sigma^{z}_{j}+h_{j}^{x}\sigma^{z}_{i}\sigma^{y}_{j})-2i\sum_{i<j<k}K_{ijk}(h_{i}^{x}\sigma^{y}_{i}\sigma^{z}_{j}\sigma^{z}_{k}+h_{j}^{x}\sigma^{z}_{i}\sigma^{y}_{j}\sigma^{z}_{k}+h_{k}^{x}\sigma^{z}_{i}\sigma^{z}_{j}\sigma^{y}_{k}).\end{split}$ (17)
Next, we compute $\Gamma_{1}(\lambda)$ as
$\Gamma_{1}(\lambda)=\text{tr}\big{[}O_{1}^{\dagger}(\lambda)O_{1}(\lambda)\big{]}=4\Big{[}\sum_{i}(h_{i}^{x}h_{i}^{z})^{2}+\sum_{i\neq j}(h_{i}^{x})^{2}J_{ij}^{2}+\sum_{i<j<k}K_{ijk}^{2}[(h_{i}^{x})^{2}+(h_{j}^{x})^{2}+(h_{k}^{x})^{2}]\Big{]}$ (18)
Following the same procedure for computing $\Gamma_{2}(\lambda)$ and finally taking the squared norm of each separated Pauli string, we obtain
$\begin{split}\Gamma_{2}(\lambda)&=16\sum_{i}(h_{i}^{x})^{2}\Big{[}(1-\lambda)h^{b}_{i}h_{i}^{z}+\lambda\Big{(}(h_{i}^{z})^{2}+\sum_{j\neq i}J_{ij}^{2}+\sum_{j,k\neq i}K_{ijk}^{2}\Big{)}\Big{]}^{2}\\ &+16\sum_{i<j}(h_{i}^{x})^{2}\Big{[}(1-\lambda)J_{ij}h^{b}_{i}+2\lambda\Big{(}J_{ij}h_{i}^{z}+\sum_{i<k<j}K_{ikj}J_{ik}+\sum_{k>j}K_{ijk}J_{ik}+\sum_{k<i}K_{kij}J_{ki}\Big{)}\Big{]}^{2}\\ &+16\sum_{i<j}(h_{j}^{x})^{2}\Big{[}(1-\lambda)J_{ij}h_{j}^{b}+2\lambda\Big{(}J_{ij}h_{j}^{z}+\sum_{i<k<j}K_{ikj}J_{kj}+\sum_{k>j}K_{ijk}J_{jk}+\sum_{k<i}K_{kij}J_{kj}\Big{)}\Big{]}^{2}\\ &+16(1-\lambda)^{2}\Big{[}\sum_{i}(h_{i}^{x})^{4}(h_{i}^{z})^{2}+\sum_{i<j}[(h_{i}^{x})^{2}+(h_{j}^{x})^{2}]^{2}J_{ij}^{2}+\sum_{i<j<k}K_{ijk}^{2}[(h_{i}^{x})^{2}+(h_{j}^{x})^{2}+(h_{k}^{x})^{2}]^{2}\Big{]}\\ &+64(1-\lambda)^{2}\Big{[}\sum_{i<j}(h_{i}^{x})^{2}(h_{j}^{x})^{2}J_{ij}^{2}+\sum_{i<j<k}K_{ijk}^{2}[(h_{i}^{x})^{2}(h_{j}^{x})^{2}+(h_{i}^{x})^{2}(h_{k}^{x})^{2}+(h_{j}^{x})^{2}(h_{k}^{x})^{2}]\Big{]}\\ &\begin{split}+16\sum_{i<j<k}(h_{i}^{x})^{2}&\Big{[}(1-\lambda)K_{ijk}h^{b}_{i}+2\lambda\Big{(}J_{ij}J_{ik}+K_{ijk}h_{i}^{z}+\sum_{j<p<k}K_{ijp}K_{ipk}+\sum_{i<p<j}K_{ipj}K_{ipk}\\ &+\sum_{p<i}K_{pij}K_{pjk}+\sum_{p>k}K_{ijp}K_{ikp}\Big{)}\Big{]}^{2}\end{split}\\ &\begin{split}+16\sum_{i<j<k}(h_{j}^{x})^{2}&\Big{[}(1-\lambda)K_{ijk}h_{j}^{b}+2\lambda\Big{(}J_{jk}J_{ij}+K_{ijk}h_{j}^{z}+\sum_{j<p<k}K_{ijp}K_{jpk}+\sum_{i<p<j}K_{ipj}K_{pjk}\\ &+\sum_{p<i}K_{pij}K_{pjk}+\sum_{p>k}K_{ijp}K_{jkp}\Big{)}\Big{]}^{2}\end{split}\\ &\begin{split}+16\sum_{i<j<k}(h_{k}^{x})^{2}&\Big{[}(1-\lambda)K_{ijk}h_{k}^{b}+2\lambda\Big{(}J_{ik}J_{jk}+K_{ijk}h_{k}^{z}+\sum_{j<p<k}K_{ipk}K_{jpk}+\sum_{i<p<j}K_{ipk}K_{pjk}\\ &+\sum_{p<i}K_{pik}K_{pjk}+\sum_{p>k}K_{ikp}K_{jkp}\Big{)}\Big{]}^{2}\end{split}\\ &\begin{split}+64\lambda^{2}\sum_{i<j<k<p}&[(h_{i}^{x})^{2}(K_{ikp}J_{ij}+K_{ijp}J_{ik}+K_{ijk}J_{ip})^{2}+(h_{j}^{x})^{2}(K_{jkp}J_{ij}+K_{ijp}J_{jk}+K_{ijk}J_{jp})^{2}\\ &+(h_{k}^{x})^{2}(K_{jkp}J_{ik}+K_{ikp}J_{jk}+K_{ijk}J_{kp})^{2}+(h_{p}^{x})^{2}(K_{jkp}J_{ip}+K_{ikp}J_{jp}+K_{ijp}J_{kp})^{2}]\end{split}\\ &\begin{split}+64\lambda^{2}\sum_{i<j<k<p<q}&[(h_{i}^{x})^{2}(K_{ijk}K_{ipq}+K_{ikq}K_{ijp}+K_{ikp}K_{ijq})^{2}+(h_{j}^{x})^{2}(K_{ijk}K_{jpq}+K_{ijp}K_{jkq}+K_{ijq}K_{jkp})^{2}\\ &+(h_{k}^{x})^{2}(K_{ijk}K_{kpq}+K_{ikp}K_{jkq}+K_{ikq}K_{jkp})^{2}+(h_{p}^{x})^{2}(K_{ijp}K_{kpq}+K_{ikp}K_{jpq}+K_{ipq}K_{jkp})^{2}\\ &+(h_{q}^{x})^{2}(K_{ijq}K_{kpq}+K_{ikq}K_{jpq}+K_{ipq}K_{jkq})^{2}].\end{split}\end{split}$ (19)
Appendix B DMRG-based estimation of ground state energies
To obtain the reference ground state energy of the nearest-neighbour Hamiltonian of Eq. (12), we employ the DMRG algorithm [51], which is an optimization method using a matrix product state (MPS) as a variational ansatz. Due to entanglement area laws [84], DMRG works particularly well for one-dimensional systems described by MPSs and for local Hamiltonians. We use the implementation in the ITensor Julia library [74]. As initial states, use random MPSs with bond dimension $\chi>1$ since setting the initial bond dimension to $1$ results in higher ground state energies. Even though the exact ground state is a product state (bitstring in the computational basis), we find that the algorithm converges to an entangled state with bond dimension larger than $1$. This is indicated by non-integer values of the occupation probabilities $\braket{n_{i}}$. To obtain an estimate for the ground state energy, we adopt the following procedure:
  1. 1.
    Obtain the ground state |$\ket{\psi_{\text{DMRG}}}$ via DMRG.
  2. 2.
    For all qubits, measure the occupation probability
    $\braket{n_{i}}=\braket{\psi_{\text{DMRG}}}{n_{i}}{\psi_{\text{DMRG}}}=\frac{1-\braket{\sigma^{z}_{i}}}{2},$ (20)
    where $0\leq\braket{n_{i}}\leq 1$.
  3. 3.
    Construct an estimate of the ground-state bit string |$\ket{\psi_{\text{GS}}}$ by rounding $\braket{n_{i}}$ to the nearest integer.
  4. 4.
    Compute an estimate of the ground-state energy $\epsilon_{\text{min}}$ by applying $H_{f}\ket{\psi_{\text{GS}}}=\epsilon_{\text{min}}\ket{\psi_{\text{GS}}}$.
  5. 5.
    Choose the smallest $\epsilon_{\text{min}}$ out of the values found for 100 random initial states with bond dimension $\chi=5$.
Note that the random coupling constants are fixed, i.e. there is no averaging over disorder configurations. Therefore, we select the minimum energy obtained by sampling over 100 randomly generated initial MPSs. The ground-state-energy estimate computed in this way is lower than the energy of each |$\ket{\psi_{\text{DMRG}}}$.
The DMRG algorithm is known to be particularly efficient when the Hamiltonian has an exact MPO representation with a small bond dimension [51]. This occurs for Hamiltonians with short-range coupling terms such as considered here, or exponentially decaying ones. The convergence of DMRG, as a variational method, may however be prevented by the algorithm getting stuck in local minima. We observe this here as the state that DMRG converges to is slightly entangled rather than the product state describing a classical spin glass. For small system sizes up to 28 qubits, the DMRG result agrees with the exact one |$\ket{\psi_{\text{exact}}}$ with a maximum error $1-|\bra{\psi_{\text{exact}}}\psi_{\text{DMRG}}\rangle|^{2}\lesssim 10^{-5}$. We also verify that the product state obtained by rounding $\braket{n_{i}}$ to the nearest integer matches |$\ket{\psi_{\text{exact}}}$ in this case.
In previous DMRG studies of classical spin glasses, a quantum wavefunction annealing method was found to improve convergence [85, 86]. As an alternative to DMRG, the ground state search can be performed by imaginary time evolution. This approach was recently used for solving QUBO problems via projected entangled pair states with flexible geometry [38].
Appendix C Circuit decomposition
An important aspect of preparing and running quantum circuits on hardware is to transpile the required quantum operations according to the corresponding native gate sets provided by the platform. These typically consist of an universal gate set containing several one-qubit gates and a single two-qubit entangling gate. For ibm_fez, its native gate set is composed by
$X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\text{ }\sqrt{X}=\frac{1}{2}\begin{bmatrix}1+i&1-i\\ 1-i&1+i\end{bmatrix},\text{ }\text{RZ}(\theta)=e^{-i\frac{\theta}{2}\sigma^{z}},$ (21)
with the entangling gate $\text{CZ}=\operatorname{diag}(1,1,1,-1)$.
The constrained connectivity among qubits of ibm_fez, and quantum platforms with heavy-hex coupling maps in general, suggests to work with highly sparse problems. The linearity of the Hamiltonians considered naturally fits into the heavy-hex IBM qubit coupling maps, whose corresponding circuit decomposition can be seen in Fig. 5.
{quantikz}
[…]
Figure 5: Circuit decomposition of a five-qubit NN HUBO instance [Eq. (1)]. The first one-qubit operator layer prepares the initial ground state |$\ket{\tilde{\psi}}=\bigotimes_{i=1}^{N}R_{y}(\theta_{i})\ket{0}_{i}$. Afterwards, $n_{\text{trot}}$ Trotter steps are applied to implement the digitized time-evolution operator of Eq. (7). Finally, all qubits are measured, and their energy distribution is used for updating the bias field of the subsequent iteration.

### v2

@@ Conclusions
In this study, we investigate and evaluate the performance of variants of the recently proposed BF-DCQO algorithm [51] by solving large-scale HUBO instances on IBM quantum processors and IonQ noisy emulators. The superiority of our BF-DCQO variant over well-established classical techniques and optimization on the D-Wave Advantage_system4.1 is verified through computing the approximation ratio and the distance to solution in an iterative process. We obtain better optimal solutions even while the circuits after each iteration demand less resources. This is achieved by means of an iterative learning procedure, where the optimal solutions obtained from the last iteration are used for the subsequent one, providing several techniques to guarantee convergence under different scenarios. Furthermore, in order to address not only the current but also the near-term hardware, we examine the expected performance of BF-DCQO on the upcoming IBM Osprey platform with 433 qubits. Simulating 433-qubit HUBO instances without noise yields promising results. The BF-DCQO algorithm does not require any classical optimization subroutines, thus avoiding potential trainability drawbacks, nor does it require extra qubits for mapping the initial HUBO into a QUBO problem.
We provide numerical evidence of how our protocol disentangles the states obtained after each iteration, one of the key signatures of our method. This feature enables an iterative reduction in the number of resources required, thus making subsequent circuits more classically tractable and opening the door to employing hybrid quantum-classical strategies. Therefore, the BF-DCQO algorithm emerges as a suitable algorithm for solving higher-order binary optimization problems on both current and near-term quantum platforms, paving the path towards solving large-scale optimization problems with status-quo and next-generation quantum processors.
Acknowledgments
@@
We would like to thank Shubham Kumar for preparing the schematic figure. We also thank Gaurav Dev and Bhargava A. Balaganchi for fruitful discussions. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team.
Data availability. Data is available from the corresponding author upon reasonable request.
Code availability. The codes used to generate data for this paper are available from the corresponding author upon reasonable request.
Competing interests. The authors declare no competing interests.
Author contributions. E.S. and N.N.H. conceived and supervised the work. S.V.R., A.-M.V., A.G.C. and N.N.H. developed the theoretical framework. S.V.R., A.-M.V., A.G.C. and A.S. developed the codes for obtaining the benchmark and hardware results. S.V.R. and N.N.H. analyzed the results. All authors participated in the discussion of results and writing the manuscript.